OpenStudy (anonymous):

In solving the equation (x + 4)(x – 7) = -18, Eric stated that the solution would be x + 4 = -18 => x = -22 or (x – 7) = -18 => x = -11 However, at least one of these solutions fails to work when substituted back into the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning

OpenStudy (anonymous):

this is the soltuion: \[ (x + 4)(x – 7) = -18\] expand left hand side: \[ (x^2-3x-28 = -18\] bring the -18 in: \[ x^2-3x-10=0\] now factor again: \[ (x-5)(x+2)=0\] so now: \[ (x-5)=0\] OR \[ (x+2)=0\] So: x=+5 OR x= -2